OpenStudy (anonymous):

I need some help solving for intersection points. I will type equation using equation editor.

OpenStudy (anonymous):

\[\sin (\frac{ \pi x }{ 2 })=x^8\]

OpenStudy (madhu.mukherjee.946):

is the answer given?

OpenStudy (anonymous):

No, but wolfram says the answers are 0 and 1.

OpenStudy (madhu.mukherjee.946):

yeah i got it

OpenStudy (madhu.mukherjee.946):

see whats the value of sin 90

OpenStudy (madhu.mukherjee.946):

@cfehr003

OpenStudy (madhu.mukherjee.946):